NAOJ GW Elog Logbook 3.2

Yuhang and Yuefan
According to the experience of Eleonora and Matteo, the yaw of PR has a very narrow range. Also, we moved -2000 counts of PR and we observed a stronger vibration of green beam. So we guess maybe PR is touching the intermedium mass.
So we checked how much we moved and them tried to move it little by little and then have a look the spectrum to see if there is any difference of peak values. The measurement was down from -1000 to 2000 counts while the original position is 750 counts. The spectrum is attached in the sequence from -1000 to 2000.
We could see that after 1300, the high frequency noise spectrum is a bit higher than the reference. And we could see at -200 and -1000, the spectrums have parts lower than the reference.

Yuefan and Yuhang
Today we finished all the cabling works for the AA telescope.
On the top left of the first picture shows the DGS board, 16 channels on the bottom are connected to the I and Q DC output of the demodulation board for both quadrants. On the top, the middle 8 channels are the DC output from the QPDio-base for two quadrants.
In the right middle of this picture, there are the two boards to control the galvo, four inputs on both boards are also from the QPDio-base DC output. and the x and y output will be connected to the galvo.
Second picture shows the demodulation board for two quadrants, one the left is the 16 demodulation signal mentioned before, and on the right, there are 8 channels which connect to the RF output of the quadrants. We didn't connect the LO yet.
Third picture shows the QPDio-base, the left four channels are the DC output, and the right two cables will be connected to the quadrant.
In the fourth picture, these are the cables that will connect the galvo motors and the RF output of the quadrants.
We put lables for all the cables we connected today to avoid any confusion in the future. We still need to organize a bit the cables, and we will also check if all of them are working tomorrow.
We also prepared the rest lens and mirrors for the telescope, if the green is availabe tomorrow we can continue the installation.

I swapped the lens to compose a cat's eye retroreflector for AOM path.
Next step is adjust the configuration and alignment by maximizing the efficiency with AOM modulation.

Activity of 09/06:
We removed spare ETMY substrate and applied first contact on the surface we didn't clean before.
We checked the calibration (figure 1 and 2) before installing TAMA size sapphire and measure the absorption.
R_surf = AC_surfref/(DC_surfref*P_in*abs_surfref) = 16.9 [1/W]
where AC_surfref = 0.40V, DC_surfref = 3.85V, P_in = 0.028W and abs_surfref = 0.22
R_bulk = AC_bulkref/(DC_bulkref*sqrt(T_bulkref)*P_in*abs_bulkref) = 0.797 [cm/W]
where AC_bulkref = 0.08V, DC_bulkref = 4.65V, T_bulkref = 0.55, P_in = 0.028W and abs_bulkref = 1.04/cm
We then moved the translation stage of the imaging unit to take into account the thickness of the TAMA size sample (60mm) from 70mm to 44.3mm.
Reminder: dL_IU = (n-1)*L_sample / n
After calibration and installation we checked the XY limits of the translation stage and modified the X one to take into account the size difference with respect to KAGRA size mirrors (IMPORTANT: change them to go back to KAGRA size).
The coordinates of the center of the sample are: [X_c, Y_c] = [327.75, 123.25]
Then we did a series of z scan to determine the z position of input and output surface. Fig 3 shows the last one performed with 10.1W of input power.
According to the phase profile the input and output surface are at Z = [44, 79]

Yuefan and Yuhang
According to Yuhang, the beam splitter after the Faraday, which is also at the bottom of the periscope was changed later last week after all the mirrors were installed. The beam doesn't seem so well aligned on the mirrors anymore.
So we tried to align again everything today, and at the same time we increase the beam height to around 224.5mm, to match the height of the quadrant on the breadboard.
During the alignment, we lost the green reflection very frequently, and at the last try when we want to install the second lens on the breadboard, the reflection beam became really low. Then we checked the end camera, half of the beam are out of the screen, and we could not use only the BS pitch local control to recover the beam height, but the PR pitch is saturating with zero corrections.

We checked the cables that connect the quadrant and board, if they are long enough for us to put the board above and below the bench. We tried to put the quadrant at the edge of the breadboard which is also the furthest position from the edge of the bench. In the attached pictures, we could see it is possible to do in both way. Anyway, I also contacted Matteo, he will go to ask about the cables next week when he will be in Nikhef.

Now the Faraday is located 10~14 holes after PBS, according to the measurement of beam parameter we could predict the beam size at both ends of this Faraday isolator. The result is shown in the attached figure 1. I also checked the aperture of FI(IO-3-1064-VHP), which is 2.7mm in diameter.
In this case, the FI aperture and beam size ratio is minimum as 2.7/0.945/2 = 1.4286. We could calculate the Gaussian beam power through an aperture is P/P0 = 1-e^(-2*(1.4286)^2) = 0.98. So in our current case, even the best-aligned beam will loss 2% of the incident power. Also, I remember that we achieved almost 97% of power transmission of FI even in the case of this beam clipping issue.
Usually, we make the beam five times smaller than the aperture. But it is really difficult to have space and meet this usual requirement. However, if we move this beam one and a half hole backward, the power cut by aperture will be 0.5%. This can be realized by moving the fork of green injection mirror (the last green mirror before going inside the chamber). Because this is blocking the way of moving backward.

The reflectivity of PBS is optimized with the same method of the optimization of the dichroic mirror(HBSY11).
However, there is a strange thing, which is the reflected field is even stronger than the incident field. See attached figure 1 and 2 (incidence and reflection). So the reflection is 109/106 = 102.83%. But anyway, the reflection is maximized.

[Aritomi, Yuhang]
We measured free running and closed loop CC2 phase noise from filter cavity (attached picture). CC2 error signal is 356 mVpp. Phase noise from filter cavity is large at low frequency. Although phase noise at low frequency is suppressed by CC2, bump at low frequency in squeezing spectrum might be phase noise of CC2. Odd number harmonics of 50Hz peaks appear when CC2 is closed. These peaks are related to electronics of CC2.

[Aritomi, Yuhang]
From previous measurement, dichroic mirror we are using (HBSY11) had only 96% reflectivity while the spec reflectivity is 99.3%. We tweaked angle of the dichroic mirror and measured BAB peak height when OPO is scanned. We improved the reflection by 4% from 95.2mV to 99.2mV. However, the optimal angle is not 45deg and we lose 18% of BAB at PBS after OPO with this angle of dichroic mirror. We tried to change the angle of the PBS by hand, but it didn't improve. We need to align the PBS.

[Aritomi, Yuhang]
We measured loss from PBS after OPO to filter cavity reflection just after PR chamber. We haven't measured loss from filter cavity reflection to homodyne yet.
position | BAB power (uW) |
after PBS | 273 |
after faraday on the bench | 254 |
after HWP | 248 |
before PR chamber | 242 |
after PR chamber | 211.5 |
From this measurement, loss from PBS after OPO to filter cavity reflection is 22.5%. Since we know that loss when squeezed light is directly injected to homodyne is 21%, total loss should be at least 40%.
We found that HWP we were using before had 5% loss and seems dirty. This explains 4% additional loss when I did additional loss measurement using this HWP. After replacing with new one, HWP loss becomes 2.4%.
Faraday on the bench has 7% loss. Alignment of the faraday should be optimized.

Current squeezing level is 2.4dB. Phase noise from laser around 10kHz becomes smaller. Spectrum at low frequency is bad due to CC2 lock loss during the measurement.

[Aritomi, Yuhang]
First we maximized IR reflection from filter cavity while we couldn't find IR resonance at that point. The reflection was 84% of injection.
Then we aligned LO and BAB reflected by unlocked filter cavity into AMC and measured visibility.
Calculated visibility is somehow more than 100% but it should be good anyway. Good news is that visibility is stable although we have jittering.
Finally we measured squeezing spectrum when squeezed light is reflected by unlocked filter cavity (attached picture). Injected green is 40mW and demodulation phase for squeezing is 190deg. We have 1.8dB squeezing down to 60Hz. Since it seems we have much larger phase noise from filter cavity (It will be reported later), we may have more squeezing when pump green is lower. Large bump below 60Hz is unknown. Noise around 10kHz should be phase noise of laser since we turned it on today.
Current problems:
1. Large bump below 60Hz
2. Large phase noise from filter cavity and CC2 control is not so stable

I aligned reflection from filter cavity to AMC to see jittering effect while IR injection is still not aligned. Attached picture shows BAB mode matching in AMC when alignment is good. Largest peak is TEM00 and small peak next to TEM00 is pitch misalignment. We can see pitch jittering effect from attached movie.

[Aritomi, Yuhang, Yuefan (remotely)]
Yesterday we found IR reflection from filter cavity was only 33% of injection and IR on PR reference was 15% of injection. The reason was that polarization after faraday on the bench we recently installed was not s-polarization due to faraday rotator and some of BAB were lost in PBS before and after faraday in PR chamber. So we put HWP just after faraday on the bench and optimized the polarization. As a result, reflection from filter cavity becomes 85%.

Here is the measuremrnt of green reflection jittering at 1pm, 6pm and 9pm.
It seems that they are similar.

Aritomi and Yuhang
Today we realigned again the GR.
GR on the PR chamber target was fine. (See attached figure 1)
GR on the BS target was moved in the pitch direction. If I remember well, Eleonora intentionally removes the DC offset of PR local control. And after about two weeks, the pitch was misaligned. Maybe this means PR mirror has pitch drift without local control. (See attached figure 2)
Then we moved picomotor of PR to recover GR on the BS target.
We moved picomotor of BS to make beam go through the first iris. We also moved picomotor of BS again and made beam go through the second iris. After that, we found the transmission was not hitting the center of the camera. Since we think two irises along the filter cavity should be the best reference, we decided to move the camera to center beam on the camera. But we didn't move the fork of camera, so if this should not be done we can also recover easily. The situation before we moved camera was attached figure 3.

Aritomi and Yuhang
We found that the incident beam into the filter cavity actually it is not jittering.
But we confirm that ITM and ETM are jitter. For ITM, it is obvious that reflection is jittering. For ETM, we found ETM reflection and it was moving.
Please check the attached video. This is the green incident beam on the second target. https://drive.google.com/open?id=1O6HZabKGtyzbuyYSPC_QGHkkhuE3W1YI

Simon, Matteo
We decided to focus on some more preciser maps on spare ETMY, especially in YZ, since in the full map XZ we could see some layer-like stuctures but not in the full YZ map, which should be the case. This is probably due to a lower contrast as in YZ we have a much larger range in the absorption coefficent.
The results of the measurements can be shown in the attached figures. We can indeed see now those layer-structures in both XZ and YZ planes, as we expected.
About the reason of these layers, we can only speculate but it seems that they have their origin in some systematic oscillations of impurity concentrations during the crystallization process.
In addition to that, I finished writing the 3D representation part of the absorption maps in Python by using Myavi. I attached also a picture of the results from the last full-map measurements.

Aritomi and Yuhang
After the simulation, we implemented the filter cavity reflection telescope. It was quite strange that I moved a lot to match this beam into AMC. Maybe this is because of the injection is not well mode matched, so the reflection is also not in good shape. But anyway, we put this telescope and tried a lot to improve the matching.
The preliminary result is as the attached figures.
In the first attached figure, we could see there is a mode mismatch peak with a height of 540mV.
In the second attached figure, we could see there is a TEM00 with a height of 6.6V.
In the third and fourth figure, you can find a peak just beside the mode-mismatch peak. It changes height because of the beam jittering in pitch direction. And it can have height up to 860mV.
In the fifth and sixth figure, you can see the TEM00 peak can also change from 5.5V to 7.5V.
Conclusion:
1. We have roughly mode mistach now for filter cavity into AMC as ~7.5%.
2. Beam jittering brings misalignment of up to ~12%.
3. The estimation is not precise because the total reflected IR power also fluctuates. From the experience of green reflection, we see much more stable green reflection after using smaller focal length lens. Maybe we are having this beam clipping issue also for this measurement in IR reflection. We should check this tomorrow.
Found on 5th September, the beam was clipped!